h/2 - h/h-4= -4/h-4
answer is h does not = 4, h = 2. is this correct?

Answer by checkley71(8403)

H/2-H/(H-4)=-4/(H-4)
H/2=H/(H-4)-4/(H-4)
H/2=(H-4)/(H-4)
H/2=1 CROSS MULTIPLY
H=2*1
H=2 ANSWER.